OpenStudy (anonymous):

What do the results of this synthetic division mean? How do I get it in to this form?

OpenStudy (anonymous):

I want to \[\frac{ x^2 +2x }{ (x+1)^2 } = 1-\frac{ 1 }{ x^2+2x+1 }\]

OpenStudy (anonymous):

I got s+1 from dividing x+1 with x^2 +2x

OpenStudy (anonymous):

do I have to do it twice since my denominator is (x+1)^2?

OpenStudy (anonymous):

and I also got a remainder of -1

OpenStudy (anonymous):